How many quarters do you need to make a 10.0 ft high pile of quarters? The thickness of a quarter is 1.75 mm. [1 ft = 0.3048 m]
Answer #1

10ft high pile of quarters

10*0.3048m   = 3.048m

                      = 3.048*1000mm

                        = 3048mm

The thickness of quarter = 1.75mm

no of quarters   = 3048/1.75    = 1741.7 >>>>answer